# A new expansion of the Riemann zeta function

### v1

@@ 3 An expansion of Zeta in the critical strip
As we have seen in the Introduction the polynomials $P_{m}$ are related to the Laguerre polynomials by Mellin transform. We now give another approach to these polynomials.
@@ 3.1 The sequence of polynomials $(P_{m})$
We start with the observation that for $s\in\mathbb{C}$, and for every $n>0$, we have
$\frac{1}{n^{s}}=\frac{1}{n+1}G_{s}(\frac{n-1}{n+1}),$ (18)
where $G_{s}$ is the function defined on $D(0,1)=\{z\in\mathbb{C};|z|<1\}$ by
$G_{s}:z\mapsto\frac{2}{1-z}\Big({\frac{1-z}{1+z}}\Big)^{s}$ (19)
Since $\Re\big({\frac{1-z}{1+z}}\big)>0$ for $z\in D(0,1)$, we see that the function $G_{s}$ is analytic on $D(0,1)$. Let $P_{m}(s)$ denote the coefficients of its Taylor expansion
$\frac{2}{1-z}\Big({\frac{1-z}{1+z}}\Big)^{s}=\sum_{m=0}^{+\infty}P_{m}(s)z^{m}\ \text{ for }\ \ \ z\in D(0,1)$ (20)
Let $z=\frac{x}{1+x}$, with $x\in]0,1[$, then, using (20), we get
$\sum_{m=0}^{+\infty}P_{m}(s)\Big(\frac{x}{1+x}\Big)^{m+1}=2x(1+2x)^{-s}=\sum_{k=0}^{+\infty}(-1)^{k}2^{k+1}\frac{(s)_{k}}{k!}x^{k+1},$ (21)
where $(s)_{k}=\frac{\Gamma(s+k)}{\Gamma(s)}=s(s+1)\cdots(s+k-1)$ if $k\geq 1$ and $(s)_{0}=1.$ We use now the well-known equivalence related to Euler summation (cf.[H1])
@@
which gives, for $P_{m}(s),$ the following explicit expression
@@
We also have an integral expression for $P_{m}(s)$ when $0<\Re(s)<1$, since for $k\geq 0$, we can write
@@
Now if $0<\alpha<1$ we have, for every $x>0$
$\int_{0}^{1}{(1-u)^{\alpha-1}}u^{x}du=\int_{0}^{+\infty}{e^{-t}}{(1-e^{-t})^{\alpha-1}}e^{-xt}dt\leq\int_{0}^{+\infty}t^{\alpha-1}e^{-xt}dt=\Gamma(\alpha)x^{-\alpha},$
(the last inequality is justified by $1\leq\frac{t}{1-e^{-t}}\leq{e^{t}}$, for every $t>0$). Finally we obtain
@@
It is easily verified (cf. [G-R]) that, for $m\geq 1$, we have $P_{m}(s)=2F(-m,s,1,2),$ where $F$ is the Gauss hypergeometric function
$F(a,b,c,z)=\sum_{k\geq 0}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\frac{z^{k}}{k!},$
@@
3.2 Orthogonality relations
Let $(\varphi_{m})_{m\geq 0}$ the sequence of functions
$\varphi_{m}:x\mapsto\sqrt{2}\ e^{-x}L_{m}(2x).$
It is well known (cf. [Leb]) that the sequence $(\varphi_{m})_{m\geq 0}$ is an orthonormal system in $(L^{2}(0,+\infty),dx)$. We verify as in the introduction that, by Mellin transform, we have for $\Re(s)>0$
Thus the Mellin-Parseval formula (cf. [T]), gives
with $\delta$ the usual Kronecker symbol. That is
Since
then
This proves that the sequence $(Q_{m})_{m\geq 0}$ of polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k}\ \frac{(\frac{1}{2}+it)_{k}}{k!}$ (27)
is an orthonormal system in $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$. It is also an orthonormal basis of this space since for $|a|<\pi$, we have
$\int_{-\infty}^{+\infty}e^{a|t|}\frac{dt}{\cosh(\pi t)}<+\infty$
By a classical result on orthogonal polynomials we deduce that, for every $m\geq 1$, the roots of $Q_{m}$ are real. This implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
For the first values of $m$ we get
$Q_{0}(t)$ $=$ $1,$
$Q_{1}(t)$ $=$ $i(-2t),$
$Q_{2}(t)$ $=$ $-2t^{2}+\frac{1}{2},$
$Q_{3}(t)$ $=$ $i(\frac{4}{3}t^{3}-\frac{5}{3}t),$
$Q_{4}(t)$ $=$ $\frac{2}{3}t^{4}-\frac{7}{3}t^{2}+\frac{3}{8}.$
Remark 7.
By (24) and (25) we have, for every $m\geq 0$, the relations
$\overline{Q_{m}(t)}=Q_{m}(-t)=(-1)^{m}Q_{m}(t),$
$\big(\frac{1}{2}+it\big)Q_{m}(t-i)+\big(\frac{1}{2}-it\big)Q_{m}(t+i)=(2m+1)Q_{m}(t).$
@@ 3.3 Expansion of $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$
We now apply Ramanujan’s summation on both sides of (18). Using (20), we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}G_{s}(\frac{n-1}{n+1})=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$
The following theorem says that, for $0<\Re(s)<1$, these last two sums can be inverted.
@@
For $0<\Re(s)<1$ we have
@@
By definition 2, we can write
@@
with $c_{0}=1/4$ if $m=0$ and $c_{m}=0$ if $m\geq 1$. We will now prove that the series
$\sigma(s)=\sum_{m\geq 0}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$
is convergent for every $0<\Re(s)<1$, and that its sum is equal to
Since $c_{0}P_{0}(s)=1/2$, then, by definition 2, the series $\sigma(s)$ is also
$\sigma(s)=\frac{1}{2}+i\sum_{m\geq 0}\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt\ P_{m}(s).$ (29)
@@
is convergent and
$\int_{0}^{+\infty}\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \frac{1}{e^{2\pi t}-1}dt<+\infty.$ (31)
Proof of the lemma.
@@
End of the proof of the lemma.
By the preceding lemma the series $\sigma(s)$ is convergent for every $0<\Re(s)<1$, and we get
Using (20) with $z=\frac{\pm it}{2\pm it}$, this last expression is equal to
$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\big(\frac{1}{1+it}\big)^{s}-\big(\frac{1}{1-it}\big)^{s}\Big)\frac{1}{e^{2\pi t}-1}\ dt,$
which is by definition 2, also equal to
$\frac{1}{2}+\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}-\frac{1}{2}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}.$
The relation (28) follows immediately from (17).
3.4 Evaluation of the sum $z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}$
If we apply the shift formula of corollary 1 to the function $f:x\mapsto\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}$, we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}+y_{m}$
with
$y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}dx=\int_{1}^{2}\Big(1-\frac{2}{x}\Big)^{m}\frac{1}{x}dx.$
Now the binomial expansion
@@ 3.4 Evaluation of the sum $z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}$
gives
$z_{m}=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k+1}}-(-1)^{m}+\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\int_{1}^{2}\frac{1}{x^{k+1}}dx.$
Thus we obtain $z_{m}=x_{m}+y_{m}$, where $x_{0}=\gamma-1,\ \ y_{0}=\log(2),$ and for $m\geq 1$
$y_{m}$ $=$ $\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\frac{1}{k}\Big(1-\frac{1}{2^{k}}\Big).$
For the first values of $m$ we get
$x_{0}$ $=$ $\gamma-1,$
$x_{1}$ $=$ $\gamma-2\zeta(2)+3,$
$x_{2}$ $=$ $\gamma-4\zeta(2)+4\zeta(3)+1,$
$x_{3}$ $=$ $\gamma-6\zeta(2)+12\zeta(3)-8\zeta(4)+11/3,$
$x_{4}$ $=$ $\gamma-8\zeta(2)+24\zeta(3)-32\zeta(4)+16\zeta(5)+5/3.$
@@
This follows from the expansion (28) of theorem 6, and the combinatorial identity
@@ 3.5 An expansion of Zeta
$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\sum_{m=0}^{+\infty}\Big(\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}dx\Big)P_{m}(s)=\int_{0}^{1}\sum_{m=0}^{+\infty}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)\frac{1}{x+1}dx.$ (35)
@@
and we immediately check, with the lemma 1, that the latter series is convergent.
The relations (18) and (20) allows us to write for every $x>0$
$\sum_{m=0}^{+\infty}\frac{1}{x+1}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)=\frac{1}{x^{s}},$
and by (35) we have
$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\int_{0}^{1}\frac{1}{x^{s}}dx=-\frac{1}{s-1}.$

### v2

@@ 3.1 The sequence of polynomials $(P_{m})$
For $m\geq 0$}, let $\psi_{m}$ the function defined in the half-plane $\{Re(z)>0\}$ by
$\psi_{m}(z)=\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m},$ (18)
and let $(P_{m}(s))$ is the sequence of polynomials given by the generating function
$\frac{2}{1-\xi}\Big({\frac{1-\xi}{1+\xi}}\Big)^{s}=\sum_{m=0}^{+\infty}P_{m}(s)\ \xi^{m}\ \text{ for }\ \ \ \xi\in D(0,1).$ (19)
Then, with $\xi=\frac{z-1}{z+1}$, we have for $s\in\mathbb{C}$ and $Re(z)>0$, the expansion
$z^{-s}=\sum_{m=0}^{+\infty}\psi_{m}(z)\,P_{m}(s).$ (20)
To obtain an explicit expression of $P_{m}(s)$, let $\xi=\frac{x}{1+x}$, with $x\in]0,1[$, then by (19) we get
$\sum_{m=0}^{+\infty}P_{m}(s)\Big(\frac{x}{1+x}\Big)^{m+1}=2x(1+2x)^{-s}=\sum_{k=0}^{+\infty}(-1)^{k}2^{k+1}\frac{(s)_{k}}{k!}x^{k+1},$
where $(s)_{k}=\frac{\Gamma(s+k)}{\Gamma(s)}=s\cdots(s+k-1)$ if $k\geq 1$ and $(s)_{0}=1.$ Now the well-known equivalence (cf.[H1])
@@
gives for $P_{m}(s)$ the following expression
@@
It is easily verified that $P_{m}(s)$ is related, for $\Re(s)>0,$ with the Laguerre polynomial $L_{m}$ by
$P_{m}(s)=\frac{2}{\Gamma(s)}\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx.$ (22)
We have another integral expression for $P_{m}(s)$ when $0<\Re(s)<1$, since for $k\geq 0$, we can write
@@
Now if $0<\alpha<1$ we have
$\int_{0}^{1}{(1-u)^{\alpha-1}}u^{m}\ du=\int_{0}^{+\infty}{e^{-t}}{(1-e^{-t})^{\alpha-1}}e^{-mt}\ dt\leq\int_{0}^{+\infty}t^{\alpha-1}e^{-mt}\ dt=\Gamma(\alpha)\ m^{-\alpha},$
(the last inequality is justified by $e^{-t}\leq\frac{1-e^{-t}}{t}\leq{1}$, for every $t>0$). With $\alpha=a$ and $\alpha=1-a$, we obtain
@@
The polynomials $P_{m}$ are related with Gauss hypergeometric function
$F(a,b,c,z)=\sum_{k\geq 0}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\frac{z^{k}}{k!}.$
It is easily verified with (21) that, for $m\geq 1$, we have
$P_{m}(s)=2F(-m,s,1,2).$
@@ 3.2 Expansion of $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$
For $0<\Re(s)<1,$ we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}z_{m}P_{m}(s)\ \text{ where }\ z_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}.$ (26)
@@
We apply (20) with $z=n$ integer $\geq 1$, this gives
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\psi_{m}(n)P_{m}(s)=\sum_{n\geq 1}^{\mathcal{R}}\sum_{m=0}^{+\infty}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s).$ (28)
We now prove that, for $0<\Re(s)<1$, these last two sums can be inverted, that is to say, the series
$\sum_{m\geq 0}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}P_{m}(s)$ (29)
is convergent for every $0<\Re(s)<1$, and its sum is equal to
By definition (6), we can write
@@
with $c_{0}=1/4$ if $m=0$ and $c_{m}=0$ if $m\geq 1$. Since $c_{0}P_{0}(s)=1/2$, then the series (29) is
$\frac{1}{2}+i\sum_{m\geq 0}\int_{0}^{+\infty}\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\frac{1}{e^{2\pi t}-1}\ dt\ P_{m}(s).$
@@
is convergent for every $t>0$, and its sum verify
$\int_{0}^{+\infty}\sum_{m=0}^{+\infty}\Big|\Big(\frac{1}{2+it}\big(\frac{it}{2+it}\big)^{m}-\frac{1}{2-it}\big(\frac{-it}{2-it}\big)^{m}\Big)\ P_{m}(s)\Big|\ \frac{1}{e^{2\pi t}-1}\ dt<+\infty.$
By this lemma the series (29) is convergent for every $0<\Re(s)<1$, and we get
Using (19) with $\xi=\frac{\pm it}{2\pm it}$, this last expression is equal to
$\frac{1}{2}+i\int_{0}^{+\infty}\Big(\big(\frac{1}{1+it}\big)^{s}-\big(\frac{1}{1-it}\big)^{s}\Big)\frac{1}{e^{2\pi t}-1}\ dt=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}.$
The relation (27) follows immediately from (17).
Proof of lemma 2.
@@
3.3 Evaluation of the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}$
Lemma 3.
For any integer $m\geq 0$
$z_{m}:=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=x_{m}+y_{m}$
with $x_{0}=\gamma-1,\ \ y_{0}=\log(2),$ and for $m\geq 1$
and
$y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ \ dx=\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\frac{1}{k}\Big(1-\frac{1}{2^{k}}\Big).$ (32)
Proof.
We apply the shift formula of corollary 1 to the function $f:x\mapsto\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}$, then we get
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}+\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ dx$
For any $m$ integer $\geq 0$, let
$x_{m}=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n}\Big(1-\frac{2}{n}\Big)^{m}-(-1)^{m}\ \text{ and }\ y_{m}=\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ dx$
We have $x_{0}=\gamma-1,\ \ y_{0}=\log(2).$ And by the binomial expansion
@@ 3.3 Evaluation of the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n+1}\Big(\frac{n-1}{n+1}\Big)^{m}$
we get for $m\geq 1$
$x_{m}$ $=$ $\gamma+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k}}-(-1)^{m}.$
and
$y_{m}=\int_{1}^{2}\frac{1}{x}\big(1-\frac{2}{x}\big)^{m}\ \ dx=\sum_{k=0}^{m}\binom{m}{k}(-2)^{k}\int_{1}^{2}\frac{1}{x^{k+1}}\ dx.$
The conclusion follows from (17) which gives for $k\geq 1$
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{k}}=\zeta(k+1)-\frac{1}{k}.$
3.4 An expansion of Zeta
@@
This follows from the expansion (27) of theorem 6. Since by lemma 3, we have
$z_{m}=x_{m}+y_{m}=\gamma+\log(2)+\sum_{k=1}^{m}\binom{m}{k}(-2)^{k}\zeta(k+1)+(-1)^{m+1}+\sum_{k=1}^{m}\binom{m}{k}\frac{(-1)^{k-1}}{k},$
and we can use the combinatorial identity
@@
$\sum_{m=0}^{+\infty}y_{m}P_{m}(s)=\sum_{m=0}^{+\infty}\Big(\int_{0}^{1}\Big(\frac{x-1}{x+1}\Big)^{m}\frac{1}{x+1}\ \ dx\Big)P_{m}(s)=\int_{0}^{1}\sum_{m=0}^{+\infty}\frac{1}{x+1}\Big(\frac{x-1}{x+1}\Big)^{m}P_{m}(s)\ dx.$ (35)
@@
and we immediately check, with lemma 1, that the latter series is convergent.
The relation (34) is then a consequence of (20) and (35).
@@
For the first values of $m$ we get
$x_{1}$ $=$ $\gamma-2\zeta(2)+3,$
$x_{2}$ $=$ $\gamma-4\zeta(2)+4\zeta(3)+1,$
$x_{3}$ $=$ $\gamma-6\zeta(2)+12\zeta(3)-8\zeta(4)+11/3$
$x_{4}$ $=$ $\gamma-8\zeta(2)+24\zeta(3)-32\zeta(4)+16\zeta(5)+5/3.$
3.5 Orthogonality relations
The polynomial $P_{m}$ is related to the Mellin transform of the function
$\varphi_{m}:x\mapsto\sqrt{2}\ e^{-x}L_{m}(2x),$ (38)
since we verify easily with (22) that, for $\Re(s)>0$, we have
It is well known (cf. [Leb]) that the sequence $(\varphi_{m})_{m\geq 0}$ is an orthonormal system in $(L^{2}(0,+\infty),\ dx)$.
Theorem 9.
The sequence $(Q_{m})_{m\geq 0}$ of polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$ (40)
is an orthonormal system in $\big(L^{2}(\mathbb{R}),\frac{dt}{\cosh(\pi t)}\big)$. This implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
Proof.
By the Mellin-Parseval formula (cf. [T]), we have
with $\delta$ the usual Kronecker symbol. That is
Since
then
By a classical result on orthogonal polynomials we deduce that, for every $m\geq 1$, the roots of $Q_{m}$ are real. Thus the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$. ∎
Remark 7.
By (24) we have, for every $m\geq 0$, the relations
$\overline{Q_{m}(t)}=Q_{m}(-t)=(-1)^{m}Q_{m}(t),$
thus we have for $t\in\mathbb{R}$
$\Re\zeta(\frac{1}{2}+it)$ $=$ $2\sum_{m=0}^{+\infty}x_{2m}Q_{2m}(t)$
$\Im\zeta(\frac{1}{2}+it)$ $=$ $2\sum_{m=0}^{+\infty}x_{2m+1}Q_{2m+1}(t).$
By (25) we have,
$TQ_{m}=(2m+1)Q_{m},$
where $T$ is the operator on $\mathbb{R}[t]$ defined by
$Tf(t)=\big(\frac{1}{2}+it\big)f(t-i)+\big(\frac{1}{2}-it\big)f(t+i).$